OpenStudy (anonymous):

Find the polynomial of least degree, having integral coefficients and leading coefficients equal to 1, with square root of 3 and square root of 2 as a zero.

OpenStudy (debbieg):

Each root c has to have an associated factor, x-c. Since \(\sqrt{3}\) is a root, you must have \((x-\sqrt{3})\) as a factor. However, in order to have integer coefficients, you must then also have \((x+\sqrt{3})\) as a factor. so you can start with: \((x-\sqrt{3})(x+\sqrt{3})\)

OpenStudy (debbieg):

Similiarly for \(\sqrt{2}\), you will have to have \((x-\sqrt{2})\) as a factor - and what else? Multiply it all out, and you'll have your polynomial that meets the specified criteria.

OpenStudy (anonymous):

so for the square root of 3.. I'll have an equation of x^2 - 3

OpenStudy (debbieg):

Well, you'll have a FACTOR of \((x^2-3)\), yes. (That isn't an equation. :)

OpenStudy (anonymous):

then I'll have x^2 - 2

OpenStudy (debbieg):

right... now multiply those two factors together....

OpenStudy (anonymous):

sorry.. and what next?

OpenStudy (debbieg):

\((x^2-3)(x^2-2)\)

OpenStudy (debbieg):

Just use FOIL.

OpenStudy (anonymous):

so I'll have x^4-5x^2+6

OpenStudy (anonymous):

Is that correct?

OpenStudy (debbieg):

yup! :)

OpenStudy (anonymous):

WOOOH! Thanks a lot! :D

OpenStudy (debbieg):

you're welcome. :)